# CYCle: Choosing Your Collaborators Wisely to Enhance Collaborative Fairness in Decentralized Learning

### v1

@@ 4 Proposed CYCle Protocol
Assumptions and Scope: In this work, we restrict ourselves to the PDL framework presented in Eqs. 1 and 2. We also assume a cross-silo setting, where the number of participants is relatively small ($N<10$) to ensure practical feasibility of PDL. Our method requires a PDL algorithm that can compute pairwise DL in a privacy- preserving way. While we employ CaPriDe learning [25] as the baseline method for PDL in this work, other alternatives can also be used. Complete description of the baseline PDL method and discussions about its convergence, communication efficiency, or privacy guarantees are beyond the scope of this work.
Rationale: To maximize its individual CG (and hence maximize MCG), each participant must intuitively give more importance to distillation losses corresponding to other reliable participants, while de-emphasizing distillation losses corresponding to unreliable participants. To minimize CGS, each participant must penalize other unreliable participants (who do not contribute to its own learning) by sharing less knowledge with them. To achieve both these goals, we need to solve the following two sub-problems: (i) Reputation Scoring (RS): How to efficiently evaluate the reliability/reputation $r_{(n,k)}$ of participant $\mathcal{P}_{k}$ in the context of collaborative learning of model $\mathcal{M}_{\theta_{n}^{*}}$ belonging to $\mathcal{P}_{n}$?, (ii) Adaptive Sharing (AS): How to regulate knowledge transfer among participants based on the reputation score to achieve collaborative fairness?.
4.1 Reputation Scoring
@@
Input: Hyperparameters $\alpha$, $\tau_{opt}$, $\tau_{max}$, $\lambda_{0}$
1:ϕitalic-$\phi\leftarrow$ Is Update Available from $\mathcal{P}_{k}$?
@@
Several reputation scoring methods have been proposed in the FL literature [27, 15, 29]. These reputation scores are generally used by the FL server either for client selection or for weighted aggregation. In contrast to FL, where a global model is learned by a central server, each participant builds its own local model in PDL. Hence, it is not possible to compute a single reputation score for a PDL participant. Instead, each participant has its own local estimate of the reputation of other participants. Hence, the same participant may be assigned different reputation scores by its collaborators.
The proposed reputation scoring is based on the intuition that gradient alignment between the local CE loss and pairwise DL is a strong indicator of the utility of $\mathcal{P}_{k}$ to $\mathcal{P}_{n}$. For example, if the two gradients are completely misaligned (either because the data distributions of the $\mathcal{P}_{k}$ and $\mathcal{P}_{n}$ do not have any overlap or because $\mathcal{P}_{k}$ is malicious), attempting to learn from $\mathcal{P}_{k}$ is likely to harm the learning of $\mathcal{P}_{n}$, rather than being beneficial. On the other hand, if there is perfect alignment, the two gradients reinforce each other and accelerate the model learning at $\mathcal{P}_{n}$. In fact, if KL divergence is used for pairwise DL computation, the gradients based on the local CE loss and pairwise DL will be closely aligned only when the predictions made $\mathcal{P}_{k}$ on $\mathcal{P}_{n}$’s data closely match the ground-truth labels available with $\mathcal{P}_{n}$. This is because the CE loss measures the “distance” between $\sigma_{T}(\mathcal{M}_{\theta_{n}}(\mbox{\boldmath$x$}))$ and $y$, while the DL loss measures the “distance” between $\sigma_{T}(\mathcal{M}_{\theta_{n}}(\mbox{\boldmath$x$}))$ and $\sigma_{T}(\mathcal{M}_{\theta_{k}}(\mbox{\boldmath$x$}))$. These two losses will lead to aligned gradients only when $\sigma_{T}(\mathcal{M}_{\theta_{k}}(\mbox{\boldmath$x$}))$ is close to $y$. This justifies our choice of using gradient alignment between CE loss and pairwise DL as the basis for reputation scoring.
where $\mathcal{\nabla}_{\theta_{n}}\mathcal{L}_{CE_{n}}$ and $\mathcal{\nabla}_{\theta_{n}}\mathcal{L}_{DL_{(n,k)}}$ are the gradients of cross entropy and distillation losses, respectively, with respect to the current parameters $\theta_{n}$ of participant $\mathcal{P}_{n}$, $n\in\mathcal{N}$, and $k\in\mathcal{N}\backslash\{n\}$. Note that $s_{(n,k)}\rightarrow 0$ indicates better gradient alignment. We empirically observed that when gradient alignment $s_{(n,k)}$ is below a threshold $\tau_{opt}$, collaboration with $\mathcal{P}_{k}$ is mostly beneficial to $\mathcal{P}_{n}$. This is because as long as the predictions of $\mathcal{P}_{k}$ are closer to the ground-truth at $\mathcal{P}_{n}$, learning from $\mathcal{P}_{k}$ will help $\mathcal{P}_{n}$. In contrast, when $s_{(n,k)}$ is above a threshold $\tau_{max}$, collaboration with $\mathcal{P}_{k}$ becomes harmful to $\mathcal{P}_{n}$. Based on these observations, we compute the reputation score by applying a soft clipping function to $s_{(n,k)}$. Specifically, the reputation score in the current round $\tilde{r}_{(n,k)}=h(s_{(n,k)})$, where $h$ is defined as:
$h(s)=\max\Big{(}0,\min\Big{(}1,\dfrac{s-\tau_{max}}{\tau_{opt}-\tau_{max}}\Big{)}\Big{)}$ (5)
@@
Assumption: Each $\mathcal{P}_{n}$ has $\mathcal{D}_{n}=\{\mbox{\boldmath$x$}_{j,n},y_{j,n}\}_{j=1}^{|\mathcal{D}_{n}|}$
1:$\mathcal{M}_{\theta_{n}^{0}}\leftarrow$ Local training at $\mathcal{P}_{n},\leavevmode\nobreak\ \forall\leavevmode\nobreak\ n\in\mathcal{N}$
@@
10:         $\mathcal{L}_{DL_{n}}\leftarrow\sum_{k\in\mathcal{N}\backslash\{n\}}\lambda_{(n,k)}\cdot\mathcal{L}_{DL_{(n,k)}}$
@@
To minimize variations due to stochasticity, the reputation score is updated using a momentum factor, i.e., $r_{(n,k)}^{t}=\alpha r_{(n,k)}^{(t-1)}+(1-\alpha)\tilde{r}_{(n,k)}$, where $10<\alpha<$ 1 is the momentum hyperparameter and $r_{(n,k)}^{t}$ is the final reputation score in round $t$ ($t>0$). When $t=0$, $r_{(n,k)}^{t}=\tilde{r}_{(n,k)}$. Henceforth, we drop the index $t$ for convenience. A higher reputation score $r_{(n,k)}$ implies that updates from $\mathcal{P}_{k}$ are useful for model learning at $\mathcal{P}_{n}$. Hence, $\mathcal{P}_{n}$ can directly utilize $r_{(n,k)}$ to weight ($\lambda_{(n,k)}$) the pairwise distillation loss. While it is possible to perform reputation scoring and dynamically update the weights $\lambda_{(n,k)}$ after every collaboration round, it can also be performed periodically (after a fixed number of rounds) to minimize computational costs.
@@ 4.2 Adaptive Sharing
Once the reputation scores are computed, $\mathcal{P}_{n}$ can use $r_{(n,k)}$ to choose its collaborators wisely. Specifically, $\mathcal{P}_{n}$ can decide if it needs “to share or not to share” the distillation loss $\mathcal{L}_{DL_{(k,n)}}$ with $\mathcal{P}_{k}$. To be precise, $\mathcal{P}_{n}$ computes and sends $\mathcal{L}_{DL_{(k,n)}}$ to $\mathcal{P}_{k}$ with probability $r_{(n,k)}$. This adaptive sharing mechanism plays a key role in ensuring collaborative fairness. In order to receive updates from $\mathcal{P}_{n}$ in the current collaboration round, participant $\mathcal{P}_{k}$ needs to maintain a high reputation score with $\mathcal{P}_{n}$. This in turn requires sharing of updates that are useful for model learning at $\mathcal{P}_{n}$ in the previous round.
@@
The above reputation-based adaptive sharing scheme has the following advantages: (1) It incentivizes $\mathcal{P}_{k}$ to share honest and useful updates with $\mathcal{P}_{n}$. Sharing noisy or malicious updates will hurt $\mathcal{P}_{k}$’s reputation score and hence, dampen $\mathcal{P}_{k}$’s ability to learn from $\mathcal{P}_{n}$. (2) If $\mathcal{P}_{n}$ finds that $\mathcal{P}_{k}$ has a low reputation score, it will send updates less frequently to $\mathcal{P}_{k}$, thereby saving valuable computational and communication resources. Note that in the PDL framework, privacy-preserving DL loss computation is often computationally expensive. However, there is one limitation in the proposed protocol. If $\mathcal{P}_{n}$ is malicious, it can deviate from the protocol and stop sending updates to $\mathcal{P}_{k}$ even though $r_{(n,k)}$ is high. While this can be addressed by penalizing participants (by reducing their reputation score) for not sharing updates, such an approach is not desirable because it will force participants to keep sharing updates even if they are not benefiting from the collaboration. Hence, we go for a compromise solution, where all the parties are forced to share updates after every $R$ collaboration rounds and the reputation scores are re-calibrated based on these responses.
4.3 Mean Estimation as a Toy Example for CYCle Protocol
Consider a simple scenario involving two clients, $N=2$, each aiming to estimate the mean of their data distributions ($\theta_{1}$ and $\theta_{2}$) by collaborating using the CYCle protocol. It can been shown that the probability of the estimate obtained through collaboration being better than a standalone estimate is lower bounded by $(1/8)\exp{\left(-1/4\gamma^{2}\right)}$, where $\gamma$ is the variance of the standalone estimate (see Theorem A.1). More importantly, this bound is independent of the data heterogeneity, which contrasts with FedAvg. More details about this analysis are available in the supplementary material.
@@ 5 Experiments
Table 1: Performance comparison on CIFAR-10 dataset: Validation accuracy evaluated with $N=5$ participants. The top section of the table presents the performance of our proposed framework compared to the FedAvg algorithm. The bottom part of the table compares collaboration gain and fairness of our proposed CYCle algorithm with existing works (rows 6-9). We employ MVA ($\uparrow$), MCG ($\uparrow$) and CGS ($\downarrow$) as evaluation metrics.
@@
Datasets: CIFAR-10 [15] is a dataset consisting of $60000$ images of size $32\times 32$ pixels, categorized into 10 classes with $6000$ images per class. It has $50000$ training and $10000$ test samples. CIFAR-100 dataset [15] shares similarities with the CIFAR-10 dataset, but it consists of 100 classes with 600 samples per class. The training and test sets are divided in the same way as in CIFAR-10. CIFAR-100 dataset is used with data augmentation (random rotation (up to 15 degrees), random crop, and horizontal flip). Fed-ISIC2019 [21] (from FLamby benchmark) is a multi-class dataset of dermoscopy images comprising of $23,247$ images with $8$ different melanoma classes and high label imbalance (ranging from 49% to less than 1% for class 0 to 7). The dataset is designed for $6$ clients based on the centers. Train/test split is: $(9930/2483),(3163/791),(2691/672),(1807/452)$, $(655/164),(351/88)$.
5.1 Experimental Setup
Baseline approaches: We consider several existing FL algorithms as baselines: vanilla FL based on FedAvg (FedAvg) [19], FL based on cosine gradient Shapley value (CGSV) [28], collaborative fairness in FL (CFFL) [18], and robust and fair FL (RFFL) [27]. The above methods are designed for FL with central orchestration. As mentioned earlier, the baseline VPDL method is based on CaPriDe learning [25]. We also evaluate standalone (SA) accuracy, where each participant trains its ML model on its local dataset without any collaboration with others. Furthermore, we report the results for these baseline methods on CIFAR-10 and CIFAR-100.
Figure 3: Per-participant performance comparison on CIFAR-10 dataset using custom CNN architecture from [28]: Validation accuracy evaluated with $N=5$ participants.
Participant data splitting strategies: We implement three types of strategies to split the training dataset among the participants: (i) homogeneous, (ii) heterogeneous, and (iii) imbalanced dataset sizes (denoted as ‘imbalanced’). Homogeneous setting refers to the case where each participant gets an equal number of data points per class. In contrast, the heterogeneous method assigns a varying number of data points to each participant, based on a Dirichlet($\delta$) distribution. The parameter $\delta$ reflects the degree of non-IID characteristics within the dataset, where smaller values of $\delta$ leads to a more heterogeneous setting, while larger values tend towards an IID setting. To determine the specific allocation, we sample $p\sim Dir_{N}(\delta)$ and assign a fraction $p_{n}$ of the total data samples to $\mathcal{P}_{n}$. We employ these settings for CIFAR-10 and CIFAR-100 with the following number of participants: $N=2,5,10$. In all our experiments, we utilize all the training samples (e.g. all $50000$ samples of CIFAR-10 and CIFAR-100). In the case of the imbalanced data distribution, we utilize a custom function that relies on parameters $\kappa$ and $m$, where $\kappa$ determines the proportion of data points received by each of the $m$ chosen participants. The remaining $(N-m)$ participants then share the remaining data among themselves. We explore this setting with $N=5$ participants and consider the following configurations: (i) imbalanced ($\kappa=0.8$, $m=1$), (ii) imbalanced ($\kappa=0.35$, $m=2$), and (iii) imbalanced ($\kappa=0.6$, $m=1$). For example, in the last case, one participant holds $60\%$ of the data, while the remaining $4$ participants get $10\%$ of the data each.
Implementation details: We keep the same training hyperparameters for all experiments. Cross-entropy loss is used for local model training in all methods. The PDL algorithms (VPDL and CYCle) use Kullback-Leibler (KL) divergence based on predictions as the distillation loss. The optimizer is stochastic gradient descent (SGD) with momentum. The learning rate is initially set to $0.1$ and is updated every $25$ rounds using a scheduler with a learning rate decay of $0.1$. For FL algorithms, we set the number of collaborating rounds to $t_{max}=100$ and the local update epoch to $1$. For the decentralized algorithms, we use $25$ epochs of local training before collaboration, followed by $75$ rounds of collaboration. The batch size is set to $128$ for CIFAR-10 and CIFAR-100, and $32$ for Fed-ISIC2019 dataset. Furthermore, $\alpha=0.5$, $\lambda_{0}=50$, $\tau_{opt}=0.25$ and $\tau_{max}=0.75$. We conduct our experiments on NVIDIA A100-PCIE-40GB GPUs on an internal cluster server, with each run utilizing a single GPU. The execution time for each run averages around $1.35$ hours.
Evaluation metrics: We assess the accuracy of the trained models based on their mean validation accuracy (MVA) and mean collaboration gain (MCG) across all participants. To evaluate collaborative fairness, we use the collaboration gain spread (CGS). The reputation scores are reported in a matrix form to determine the relative contribution of each client.
(a) Imbalanced (0.8, 1) setting
(b) Imbalanced (0.35, 2) setting

### v2

@@ 4 Proposed CYCle Protocol
Input: Momentum factor $\alpha$
[…]
@@
Assumptions and Scope.
In this work, we restrict ourselves to the PDL framework presented in Eqs. 1 and 2. We also assume a cross-silo setting, where the number of participants is relatively small ($N<10$) to ensure the practical feasibility of PDL. Our method requires a PDL algorithm that can compute pairwise DL in a privacy-preserving way. While we employ CaPriDe learning (Tastan & Nandakumar, 2023) as the baseline method for PDL in this work, other alternatives can also be used. Complete description of the baseline PDL method and discussions about its convergence, communication efficiency, or privacy guarantees are beyond the scope of this work.
Rationale. To maximize its individual CG (and hence maximize MCG), each participant must intuitively give more importance to distillation losses corresponding to other reliable participants, while de-emphasizing distillation losses corresponding to unreliable participants. To minimize CGS, each participant must penalize other unreliable participants (who do not contribute to their own learning) by sharing less knowledge with them. To achieve both these goals, we need to solve the following two sub-problems: (i) Reputation Scoring (RS): How to efficiently evaluate the reliability/reputation $r_{(n,k)}$ of participant $\mathcal{P}_{k}$ in the context of collaborative learning of model $\mathcal{M}_{\theta_{n}^{*}}$ belonging to $\mathcal{P}_{n}$?, (ii) Adaptive Sharing (AS): How to regulate knowledge transfer among participants based on the reputation score to achieve collaborative fairness?.
4.1 Reputation Scoring
Several reputation scoring methods have been proposed in the FL literature (Xu & Lyu, 2021; Zhang et al., 2021). These reputation scores are generally used by the FL server either for client selection or for weighted aggregation. In contrast to FL, where a global model is learned by a central server, each participant builds its own local model in PDL. Hence, it is not possible to compute a single reputation score for a PDL participant. Instead, each participant has its own local estimate of the reputation of other participants. Hence, the same participant may be assigned different reputation scores by their collaborators.
@@ 4.1 Reputation Scoring
Assumption: Each $\mathcal{P}_{n}$ has $\mathcal{D}_{n}=\{{\bm{x}}_{j,n},y_{j,n}\}_{j=1}^{|\mathcal{D}_{n}|}$
[…]
@@
[…]
@@
The proposed reputation scoring is based on the intuition that the gradient alignment between local CE loss and pairwise DL is a strong indicator of the utility of ${\mathcal{P}}_{k}$ to ${\mathcal{P}}_{n}$. For example, if the two gradients are completely misaligned (either because the data distributions of $\mathcal{P}_{k}$ and $\mathcal{P}_{n}$ do not have any overlap or because $\mathcal{P}_{k}$ is malicious), attempting to learn from $\mathcal{P}_{k}$ is likely to harm the learning of $\mathcal{P}_{n}$, rather than being beneficial. On the other hand, if there is perfect alignment, the two gradients reinforce each other and accelerate the model learning at $\mathcal{P}_{n}$. In fact, if KL divergence is used for pairwise DL computation, the gradients based on the local CE loss and pairwise DL will be closely aligned only when the predictions made by $\mathcal{P}_{k}$ on $\mathcal{P}_{n}$’s data closely match the ground-truth labels available with $\mathcal{P}_{n}$. This is because CE loss measures the “distance” between $\sigma_{T}(\mathcal{M}_{\theta_{n}}({\bm{x}}))$ and $y$, while DL loss measures the “distance” between $\sigma_{T}(\mathcal{M}_{\theta_{n}}({\bm{x}}))$ and $\sigma_{T}(\mathcal{M}_{\theta_{k}}({\bm{x}}))$. These losses will lead to aligned gradients only when $\sigma_{T}(\mathcal{M}_{\theta_{k}}({\bm{x}}))$ is close to $y$. This justifies our choice of using the gradient alignment between CE loss and pairwise DL as the basis for reputation scoring.
[…]
$h(s)=\max\left(0,\min\left(1,\dfrac{s-\tau_{max}}{\tau_{opt}-\tau_{max}}\right)\right).$ (7)
To minimize variations due to stochasticity, the reputation score is updated using a momentum factor, i.e., $r_{(n,k)}^{t}=\alpha r_{(n,k)}^{(t-1)}+(1-\alpha)\tilde{r}_{(n,k)}$, where $0<\alpha<1$ is the momentum hyperparameter and $r_{(n,k)}^{t}$ is the final reputation score in round $t$ ($t>0$). When $t=0$, $r_{(n,k)}^{t}=\tilde{r}_{(n,k)}$. Henceforth, we drop the index $t$ for convenience. A higher reputation score $r_{(n,k)}$ implies that updates from $\mathcal{P}_{k}$ are useful for model learning at $\mathcal{P}_{n}$. Hence, $\mathcal{P}_{n}$ can directly utilize $r_{(n,k)}$ to weight ($\lambda_{(n,k)}$) the pairwise distillation loss. While it is possible to perform reputation scoring and dynamically update the weights $\lambda_{(n,k)}$ after every collaboration round, it can also be performed periodically (after a fixed number of rounds) to minimize computational costs.
@@ 4.2 Adaptive Sharing
Once the reputation scores are computed, $\mathcal{P}_{n}$ can use $r_{(n,k)}$ to choose its collaborators wisely. Specifically, $\mathcal{P}_{n}$ can decide if it needs “to share or not to share” the distillation loss $\mathcal{L}_{DL_{(k,n)}}$ with $\mathcal{P}_{k}$. To be precise, $\mathcal{P}_{n}$ computes and sends $\mathcal{L}_{DL_{(k,n)}}$ to $\mathcal{P}_{k}$ with probability $r_{(n,k)}$. This adaptive sharing mechanism plays a key role in ensuring collaborative fairness. In order to receive updates from $\mathcal{P}_{n}$ in the current collaboration round, participant $\mathcal{P}_{k}$ needs to maintain a high reputation score with $\mathcal{P}_{n}$. This, in turn, requires the sharing of updates that are useful for model learning at $\mathcal{P}_{n}$ in the previous round.
@@
The above reputation-based adaptive sharing scheme has the following advantages: (1) It incentivizes $\mathcal{P}_{k}$ to share honest and useful updates with $\mathcal{P}_{n}$. Sharing noisy or malicious updates will hurt $\mathcal{P}_{k}$’s reputation score and hence, dampen $\mathcal{P}_{k}$’s ability to learn from $\mathcal{P}_{n}$. (2) If $\mathcal{P}_{n}$ finds that $\mathcal{P}_{k}$ has a low reputation score, it will send updates less frequently to $\mathcal{P}_{k}$, thereby saving valuable computational and communication resources. Note that in the PDL framework, privacy-preserving DL loss computation is often computationally expensive. However, there is one limitation in the proposed protocol. If $\mathcal{P}_{n}$ is malicious, it can deviate from the protocol and stop sending updates to $\mathcal{P}_{k}$ even though $r_{(n,k)}$ is high. While this can be addressed by penalizing participants (by reducing their reputation score) for not sharing updates, such an approach is not desirable because it will force participants to keep sharing updates even if they are not benefiting from the collaboration. Hence, we go for a compromise solution, where all the parties are forced to share updates after every $R$ collaboration rounds, and the reputation scores are recalibrated based on these responses.
4.3 Theoretical Comparison with FedAvg via Mean Estimation
To illustrate the theoretical benefit of CYCle over traditional FL approaches such as FedAvg (McMahan et al., 2017), we consider a simple mean estimation problem with two clients $(N=2)$, each minimizing $f_{k}(w)=(w-\theta_{k})^{2}$ using limited local samples. Let $\widehat{\theta}_{1}\sim\mathcal{N}(\theta_{1},\gamma^{2})$ and $\widehat{\theta}_{2}\sim\mathcal{N}(\theta_{2},\gamma^{2})$ denote the empirical estimates of each client, where $\gamma^{2}=\nu^{2}/N$ is the estimation variance.
In FedAvg, the clients compute the global model as $w=(\widehat{\theta}_{1}+\widehat{\theta}_{2})/2$. However, as shown in prior work (Cho et al., 2022), the probability that this shared model outperforms a standalone estimate is exponentially suppressed with increasing data heterogeneity:
$\mathbb{P}\left(F_{k}(w)\leq F_{k}(\widehat{\theta}_{k})\right)\leq 2\exp\left(-\dfrac{\gamma_{G}^{2}}{5\gamma^{2}}\right),\vskip-5.0pt$ (8)
where $\gamma_{G}^{2}=\left((\theta_{1}-\theta_{2})/2\right)^{2}$ quantifies heterogeneity. This shows that FedAvg performs poorly when the client distributions are dissimilar.
In contrast, CYCle uses a reputation-weighted aggregation: $w=(1-r/2)\cdot\widehat{\theta}_{1}+r/2\cdot\widehat{\theta}_{2},$ where the reputation score $r$ depends on the empirical distance between $\widehat{\theta}_{1}$ and $\widehat{\theta}_{2}$.
Theorem 4.1.
The probability that the model obtained through CYCle’s collaborative aggregation improves upon a standalone model is lower bounded as
$\mathbb{P}\left(F_{k}(w)\leq F_{k}(\widehat{\theta}_{k})\right)\geq\frac{1}{8}\exp\left(-\frac{1}{4\gamma^{2}}\right).$
Unlike FedAvg, this bound is independent of heterogeneity $\gamma_{G}^{2}$, highlighting CYCle’s robustness to client diversity. Full derivations and proofs are provided in Appendix A.
4.4 Extension to Gossip-SGD
CYCle can be integrated with decentralized optimization algorithms such as Gossip-SGD (Boyd et al., 2006; Koloskova et al., 2019) by modifying the communication mechanism through a dynamic, reputation-driven mixing matrix. In classical Gossip-SGD, each client averages model updates with neighbors using a static mixing matrix. We replace this matrix with one that evolves based on alignment scores between gradient directions, capturing the degree of collaborative benefit.
Formally, during each communication round, each node computes the pairwise cosine similarity with its neighbors’ gradients and maps the resulting score using a softmax function controlled by a hyperparameter $\beta$. The updated mixing matrix $W$ is then used to sample a binary interaction mask $S\sim\text{Bernoulli}(W)$, ensuring that communication occurs only with beneficial peers. This mechanism balances collaboration and robustness by probabilistically encouraging high-alignment links while discouraging noisy exchanges.
We detail the full algorithm and its integration into Gossip-SGD in Appendix C, where we also describe the mapping functions and sampling strategy. Empirical results on the Fed-ISIC2019 dataset confirm that this extension maintains fairness and improves utility under realistic, non-i.i.d. data conditions.
@@ 5 Experiments
Datasets. CIFAR-10 (Krizhevsky et al., 2009) is a dataset consisting of $60000$ images of size $32\times 32$ pixels, categorized into 10 classes with $6000$ images per class. It has $50000$ training and $10000$ test samples. CIFAR-100 dataset (Krizhevsky et al., 2009) shares similarities with the CIFAR-10 dataset, but it consists of 100 classes with 600 samples per class. The training and test sets are divided in the same way as in CIFAR-10. The CIFAR-100 dataset is used with data augmentation (random rotation (up to 15 degrees), random crop, and horizontal flip). Fed-ISIC2019 (Ogier du Terrail et al., 2022) (from FLamby benchmark) is a multi-class dataset of dermoscopy images comprising $23,247$ images with $8$ different melanoma classes and high label imbalance (ranging from 49% to less than 1% for class 0 to 7). The dataset is designed for $6$ clients based on the centers. Train/test split is: (9930/2483),(3163/791),(2691/672),(1807/452)(9930/2483),(3163/791),(2691/672),(1807/452), $(655/164),(351/88)$.
5.1 Experimental Setup
Baseline approaches. We consider several existing FL algorithms as baselines: vanilla FL based on FedAvg (FedAvg) (McMahan et al., 2017), FL based on cosine gradient Shapley value (CGSV) (Xu et al., 2021), collaborative fairness in FL (CFFL) (Lyu et al., 2020), and robust and fair FL (RFFL) (Xu & Lyu, 2021). The above methods are designed for FL with central orchestration. As mentioned earlier, the baseline VPDL method is based on CaPriDe learning (Tastan & Nandakumar, 2023). We also evaluate standalone (SA) accuracy, where each participant trains its ML model on its local dataset without any collaboration with others. Furthermore, we report the results for these baseline methods on CIFAR-10 and CIFAR-100. We do not compare against Cronus or similar decentralized methods, as they focus on robustness under adversarial conditions and omit fairness evaluation. Since our goal is to promote fair collaboration, we instead benchmark against fairness-aware FL methods. We leave integration with robustness-centric frameworks for future work.
Participant data splitting strategies. We implement three types of strategies to split the training dataset among the participants: (i) homogeneous, (ii) heterogeneous, and (iii) imbalanced dataset sizes (denoted as ‘imbalanced’). A homogeneous setting refers to the case where each participant gets an equal number of data points per class. In contrast, the heterogeneous method assigns a varying number of data points to each participant, based on a Dirichlet($\delta$) distribution. The parameter $\delta$ reflects the degree of non-IID characteristics within the dataset, where smaller values of $\delta$ lead to a more heterogeneous setting, while larger values tend towards an IID setting. To determine the specific allocation, we sample $p\sim Dir_{N}(\delta)$ and assign a fraction $p_{n}$ of the total data samples to $\mathcal{P}_{n}$. We employ these settings for CIFAR-10 and CIFAR-100 with the following number of participants: $N=2,5,10$. In all our experiments, we utilize all training samples (e.g., all $50000$ samples of CIFAR-10 and CIFAR-100). In the case of the imbalanced data distribution, we utilize a custom function that relies on parameters $\kappa$ and $m$, where $\kappa$ determines the proportion of data points received by each of the $m$ chosen participants. The remaining (Nm)(N-m) participants then share the remaining data among themselves. We explore this setting with $N=5$ participants and consider the following configurations: (i) imbalanced ($\kappa=0.8$, $m=1$), (ii) imbalanced ($\kappa=0.35$, $m=2$), and (iii) imbalanced ($\kappa=0.6$, $m=1$). For example, in the last case, one participant holds $60\%$ of the data, while the remaining $4$ participants get $10\%$ of the data each.
Table 1: Performance comparison on CIFAR-10 dataset: Validation accuracy evaluated with $N=5$ participants. The top section of the table presents the performance of our proposed framework compared to the FedAvg algorithm. The bottom part of the table compares the collaboration gain and fairness of our proposed CYCle algorithm with existing works (rows 6-9). We use MVA ($\uparrow$), MCG ($\uparrow$) and CGS ($\downarrow$) as eval. metrics.
@@ 5.1 Experimental Setup
Evaluation metrics. We assess the accuracy of the trained models based on their mean validation accuracy (MVA) and mean collaboration gain (MCG) across all participants. To evaluate collaborative fairness, we use the collaboration gain spread (CGS). Reputation scores are reported in a matrix form to determine the relative contribution of each client.